Suppose you have a sample of sodium weighing 11.25 g. How many atoms of sodium are present in the sample?
1
Expert's answer
2018-01-03T08:18:07-0500
m(Na)= 11.25 g
M(Na)= 23 g/mol
n(Na)= m/M = (11.25/23) mol = 0.48913 mol
N(Na atoms) = NA × n
Avogadro's number definition, the constant, 6.022 × 1023 , representing the number of atoms, molecules, or ions in one mole of a substance. Symbol: NA.
N(Na atoms)=6.022 × 1023 × 0.48913 = 2.945 × 1023 atoms.